7.) Assume a linear demand function of the form:
Qd = 120 - 5P
and a linear supply curve of the form:
Qs = -30 + 10P
Using these demand and supply functions, answer the following questions. Once you
have had a go at the questions, follow the link below to compare your answers.
a) Calculate the quantities demanded and supplied for prices from $3 - $15.
b) Plot these figures to give the demand and supply curves for the product.
c) Using simultaneous equations, calculate the equilibrium price and output.
1
Expert's answer
2020-05-25T09:46:18-0400

Linear demand function

Qd = 120-5P

Linear supply curve

Qs = -30 + 10P

a)     Quantities demanded and supplied for prices $3 - $15

b)     Demand and supply curves

c)     Equilibrium price and output

Equilibrium = Quantity Demanded (Qd) = Quantity Supplied (Qs)

Equilibrium price

(Qd = 120-5P) = (Qs = -30 + 10P)

1OP + 5P = 120 + 30

15P = 150

P = 150/15 =$10


Equilibrium Quantity (Qd)

Qd = 120-5P

Qd = 120-5(10)

Qd = 120-50

Qd = 70